OpenStudy (anonymous):

Solve: log2(6-2x)-log2x=3

OpenStudy (xapproachesinfinity):

use this property \(\Large \tt\color{purple}{loga+logb=loga.b}\)

OpenStudy (anonymous):

I don't really get it...

OpenStudy (xapproachesinfinity):

well a in your case is 2(6-2x) and b is 2x

OpenStudy (xapproachesinfinity):

Oh im sorry error lol

OpenStudy (xapproachesinfinity):

you need to use this property \(\Large \tt\color{purple}{log(a)-log(b)=log(\frac{a}{b})}\) i missed the minus sign xD

OpenStudy (xapproachesinfinity):

so \(\Large \tt\color{purple}{log[2(6-2x)]-log(2x)=3\\log[\frac{2(6-2x)}{2x}]=3\\log[\frac{6-2x}{2x}]=3}\) carry on the rest

OpenStudy (xapproachesinfinity):

correction! \(\Large \tt\color{purple}{log[\frac{6-2x}{x}]}\) i forgot to cancel that 2 from top and bottom

OpenStudy (xapproachesinfinity):

@AnnGrey

OpenStudy (anonymous):

I'm supposed to solve for x right?

OpenStudy (xapproachesinfinity):

correct! what can you do next?

OpenStudy (anonymous):

Multiply? I don't remember my math anymore...I lost it during the summer.

OpenStudy (xapproachesinfinity):

No you have to do inverse operation of log which exponential

OpenStudy (anonymous):

by any chance, is x = 6/(2+e^3)

OpenStudy (xapproachesinfinity):

check and see, do you know how to undo log?

OpenStudy (tkhunny):

What is all this? Are we wandering in the wilderness? Are these logs Base 2? log2(6-2x)-log2x=3 log2((6-2x)/x) = 3 2^3 = (6-2x)/x 8x = 6-2x 10x = 6 x = 6/10 = 3/5 Checking... log2(6-2(3/5))-log2(3/5)=3 log2(30/5 - 6/5)-log2(3/5)=3 log2(24/5)-log2(3/5)=3 log2((24/5)/(3/5))=3 log2((24/5)*(5/3))=3 log2((24/3)*(5/5))=3 log2((8)*(1))=3 log2(8)=3 Yup!

OpenStudy (xapproachesinfinity):

oh i thought it wasn't the base -_-

OpenStudy (tkhunny):

No worries. This notation is always difficult to translate.

OpenStudy (xapproachesinfinity):

^_^ thanks though for checking